OpenStudy (anonymous):

∫[(e^x)/(e^(2x)+3e^x+2),]

OpenStudy (anonymous):

lokisan u are you replying?

OpenStudy (anonymous):

The denominator can be factored as\[(e^x+1)(e^x+2)\]If you then set \[u=e^x \rightarrow du=e^x dx\]you'll end up with\[I=\int\limits_{}{}\frac{e^x}{e^{2x}+3e^x+2}dx=\int\limits_{}{}\frac{du}{(u+1)(u+2)}\]You can use partial fraction decomposition on this last expression:\[\frac{1}{(u+1)(u+2)}=\frac{A}{u+1}+\frac{B}{u+2}\]to find\[A=1, B=-1\]Your integral is then\[I=\int\limits_{}{}\frac{1}{u+1}-\frac{1}{u+2}du = \log (u+1)-\log(u+2)+c\]Substituting back,\[I=\log \frac{e^x+1}{e^x+2}+c\]

OpenStudy (anonymous):

is this the answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you. this looks so similar to what we are doing

OpenStudy (anonymous):

np - I'd appreciate another fan :P

OpenStudy (anonymous):

Whenever you see a polynomial in the denominator, try to factor it. Then try to see if you can use partial fraction decomposition.

OpenStudy (anonymous):

how do you become a fan? you are a lifesaver. My evil calculas teacher gave us a take home test even i missed the whole week. thank you and can you help with other problems?

OpenStudy (anonymous):

there should be a link next to my name - a blue line saying, "Become a fan". If it's not there, maybe try refreshing you page.

OpenStudy (anonymous):

dichalao, what the hell's up with your professor giving you that question? It's an algebraic pain.

OpenStudy (anonymous):

she's evil

OpenStudy (anonymous):

How many more questions do you have?

OpenStudy (anonymous):

I have 10 out of which i have done like 3 . 3 of them are set up only using fraction decomposition. and 2 are solving using decomposition

OpenStudy (anonymous):

You may want to spread those questions around (i.e. keep making new posts) so others can help too.

OpenStudy (anonymous):

do you know how can i type the equation easily like you did in the reply?

OpenStudy (anonymous):

There's a button below, "Equation". You can enter things in there.

OpenStudy (anonymous):

sorry but can you show the decomposition part?

OpenStudy (anonymous):

Loki do u have time to help me out on that problem?

OpenStudy (anonymous):

no in the first answer. you said a=1 b=-1

OpenStudy (anonymous):

Yes dichalao...just let me finish here and I'll go back.

OpenStudy (anonymous):

In partial fraction decomposition, you set it out like that when you have linear factors in the bottom. You then multiply both sides by (u+1)(u+2) and you'll end up with 1 on the left and what you see on the right. You need to find A and B that will make the statement true; that is, by comparing coefficients on the left and right.

OpenStudy (anonymous):

thanks